# Subsampled Ensemble Can Improve Generalization Tail Exponentially

### v3

@@ C.3 Proof for Theorem 2
$\mathbb{P}\left(\mathbb{P}_{*}\left(\mathcal{A}(z_{1}^{*},\ldots,z_{k}^{*})\in\Theta^{\delta}\right)<\frac{1-\mathcal{E}_{k,\delta}}{e}\right)$
$\leq$ $\min\left(\exp\left(-\frac{n}{2k}\cdot\left(1-\frac{2}{e}\right)(1-\mathcal{E}_{k,\delta})\right),\left(\frac{e^{2}\mathcal{E}_{k,\delta}}{e-1}\right)^{(1-1/e)\frac{n}{2k}}\right).$
@@
$\mathbb{P}\left(\mathcal{S}\cap\Theta^{\delta}=\emptyset\right)$
$\leq$ $\min\left(\exp\left(-\frac{n}{2k}\cdot\left(1-\frac{2}{e}\right)(1-\mathcal{E}_{k,\delta})\right),\left(\frac{e^{2}\mathcal{E}_{k,\delta}}{e-1}\right)^{(1-1/e)\frac{n}{2k}}\right)+\exp\left(-\frac{B_{1}(1-\mathcal{E}_{k,\delta})}{e}\right)$
$\leq$ $\min\left(\exp\left(-\dfrac{1-2/e}{1-1/e}\cdot(1-\mathcal{E}_{k,\delta})\right),\frac{e^{2}\mathcal{E}_{k,\delta}}{e-1}\right)^{(1-1/e)\frac{n}{2k}}+\exp\left(-\frac{B_{1}(1-\mathcal{E}_{k,\delta})}{e}\right).$
The conclusion then follows by setting $C_{4},C_{5},C_{6},C_{7}$ to be the appropriate constants. $\Box$
The second lemma gives bounds for the excess risk sensitivity $\bar{\eta}_{k,\delta}$ in the case of the set-valued learning algorithm (24):
Lemma 7 (Bounds of $\bar{\eta}_{k,\delta}$ for the set-valued learning algorithm (24))
is used with $\epsilon\geq 0$, it holds that
and hence
where $T_{k}$ is the tail probability defined in Theorem 2.
Proof of Lemma 7. Let $\hat{L}_{k}(\theta):=\frac{1}{k}\sum_{i=1}^{k}l(\theta,z_{i})$. Let $\theta^{*}$ be an optimal solution of (1). We have
@@
This immediately gives (40). (41) is obvious given (39) and (40). $\Box$
To prove Theorem 2, we introduce some notation. For every non-empty subset $\mathcal{W}\subseteq\Theta$, we use the following counterpart of Definition 3. Let
Proof of Theorem 2 for $\mathsf{ROVEs}$. Given the retrieved solution set $\mathcal{S}$ and the chosen $\epsilon$, the rest of Phase II of Algorithm 2 exactly performs Algorithm 3 on the restricted problem $\min_{\theta\in\mathcal{S}}\mathbb{E}\left[l(\theta,z)\right]$ to obtain $\hat{\theta}_{n}$ with the data $\bm{z}_{\lfloor n/2\rfloor+1:n}$, the set-valued learning algorithm (24), the chosen $\epsilon$ value and $k=k_{2},B=B_{2}$.
and hence by union bound we can write
$\mathbb{P}\left(\mathcal{S}\cap\Theta^{\delta}=\emptyset\right)$ has a bound from Lemma 6. We focus on the second probability.
If $T_{k_{2}}\left((\delta-\overline{\epsilon})/2\right)+T_{k_{2}}\left(\underline{\epsilon}/2\right)<1/5$, we have $T_{k_{2}}^{\mathcal{S}}\left((\delta-\overline{\epsilon})/2\right)+T_{k_{2}}^{\mathcal{S}}\left(\underline{\epsilon}/2\right)<1/5$ and subsequently $51-T_{k_{2}}^{\mathcal{S}}\left((\delta-\epsilon)/2\right)-T_{k_{2}}^{\mathcal{S}}\left(\epsilon/2\right)>4/$ 5, and hence $\bar{\eta}_{k_{2},\eta}\geq 1-T_{k_{2}}^{\mathcal{S}}\left((\delta-\epsilon)/2\right)-T_{k_{2}}^{\mathcal{S}}\left(\epsilon/2\right)>4/5$ by Lemma 7 for Phase II of $\mathsf{ROVEs}$ conditioned on $\mathcal{S}$ and $\epsilon$, therefore the bound (31) from Theorem 3 applies. Using the inequalities (39) and (40) to upper bound the $\min(1-p_{k}^{\max},p_{k}^{\max}-\bar{\eta}_{k,\delta})$ term in (31) gives
@@
$\mathbb{P}\left(L(\hat{\theta}_{n})>\min_{\theta\in\mathcal{S}}L(\theta)+\delta\big{|}\mathcal{S},\epsilon\right)$
$\leq$ $\left|\mathcal{S}\right|\left(3\min\left(e^{-2/5},C_{1}\max\left(T_{k_{2}}^{\mathcal{S}}\left(\dfrac{\underline{\epsilon}}{2}\right),T_{k_{2}}^{\mathcal{S}}\left(\dfrac{\delta-\overline{\epsilon}}{2}\right)\right)\right)^{\frac{n}{2C_{2}k_{2}}}+\exp\left(-\dfrac{B_{2}}{C_{3}}\right)\right)$
$=$ $\left|\mathcal{S}\right|\left(3\min\left(e^{-2/5},C_{1}T_{k_{2}}^{\mathcal{S}}\left(\dfrac{\min(\underline{\epsilon},\delta-\overline{\epsilon})}{2}\right)\right)^{\frac{n}{2C_{2}k_{2}}}+\exp\left(-\dfrac{B_{2}}{C_{3}}\right)\right)$
$\leq$ $\left|\mathcal{S}\right|\left(3\min\left(e^{-2/5},C_{1}T_{k_{2}}\left(\dfrac{\min(\underline{\epsilon},\delta-\overline{\epsilon})}{2}\right)\right)^{\frac{n}{2C_{2}k_{2}}}+\exp\left(-\dfrac{B_{2}}{C_{3}}\right)\right).$
@@
$\mathbb{P}\left(L(\hat{\theta}_{n})>\min_{\theta\in\mathcal{S}}L(\theta)+\delta\right)\leq B_{1}\left(3\min\left(e^{-2/5},C_{1}T_{k_{2}}\left(\dfrac{\min(\underline{\epsilon},\delta-\overline{\epsilon})}{2}\right)\right)^{\frac{n}{2C_{2}k_{2}}}+\exp\left(-\dfrac{B_{2}}{C_{3}}\right)\right).$
This leads to the desired bound (10) after the above bound is plugged into (47) and the bound (37) from Lemma 6 is applied with $k=k_{1}$. $\Box$
Proof of Theorem 2 for $\mathsf{ROVE}$. For every non-empty subset $\mathcal{W}\subseteq\Theta$ and $k_{2}$, we consider the indicator
$\epsilon_{0}$ $:=0,$
$\epsilon_{m^{\prime}+1}$ $:=\dfrac{\delta}{2},$
and that
Therefore
@@
$\sup_{\epsilon\in[0,\delta/2]}\left|\bar{p}_{k_{2}}^{\mathcal{W},\epsilon}(\theta)-p_{k_{2}}^{\mathcal{W},\epsilon}(\theta)\right|$ (48)
$\leq$ $\max_{0\leq i\leq m^{\prime}+1}\max\left(\left|\bar{p}_{k_{2}}^{\mathcal{W},\epsilon_{i}}(\theta)-p_{k_{2}}^{\mathcal{W},\epsilon_{i}}(\theta)\right|,\left|\bar{p}_{k_{2}}^{\mathcal{W},\epsilon_{i}-}(\theta)-p_{k_{2}}^{\mathcal{W},\epsilon_{i}-}(\theta)\right|\right)+\dfrac{1}{m}.$
Therefore we can write
@@
$\mathbb{E}\left[\left(\max_{0\leq i\leq m^{\prime}+1}\max\left(\left|\bar{p}_{k_{2}}^{\mathcal{W},\epsilon_{i}}(\theta)-p_{k_{2}}^{\mathcal{W},\epsilon_{i}}(\theta)\right|,\left|\bar{p}_{k_{2}}^{\mathcal{W},\epsilon_{i}-}(\theta)-p_{k_{2}}^{\mathcal{W},\epsilon_{i}-}(\theta)\right|\right)\right)^{2}\right]$
By Minkowski inequality, the supremum satisfies
Recall the bound (43) from the proof of Lemma 7. Here we have the similar bound $\max_{\theta\in\mathcal{W}\backslash\mathcal{W}^{\delta}}p_{k_{2}}^{\mathcal{W},\epsilon}(\theta)\leq\mathbb{P}\left(\widehat{\mathcal{W}}_{k_{2}}^{\epsilon}\not\subseteq\mathcal{W}^{\delta}\right)$, and hence
$\hat{L}_{k_{2}}(\theta)$ $:=\frac{1}{k_{2}}\sum_{i=1}^{k_{2}}l(\theta,z_{i}),$
and have
$\left\{\widehat{\mathcal{W}}_{k_{2}}^{\delta/2}\not\subseteq\mathcal{W}^{\delta}\right\}$
$\subseteq$ $\bigcup_{\theta,\theta^{\prime}\in\mathcal{W}\text{ s.t. }L(\theta^{\prime})-L(\theta)>\delta}\left\{\hat{L}_{k_{2}}(\theta^{\prime})\leq\hat{L}_{k_{2}}(\theta)+\dfrac{\delta}{2}\right\}$
$\subseteq$ $\bigcup_{\theta,\theta^{\prime}\in\Theta\text{ s.t. }L(\theta^{\prime})-L(\theta)>\delta}\left\{\hat{L}_{k_{2}}(\theta^{\prime})-L(\theta^{\prime})+\max(\Delta,\delta)\leq\hat{L}_{k_{2}}(\theta)-L(\theta)+\dfrac{\delta}{2}\right\}$
  by the definition of $\Delta_{o}$
$\subseteq$ $\bigcup_{\theta,\theta^{\prime}\in\Theta}\left\{\hat{L}_{k_{2}}(\theta^{\prime})-L(\theta^{\prime})+\max\left(\Delta_{o}-\dfrac{\delta}{2},\dfrac{\delta}{2}\right)\leq\hat{L}_{k_{2}}(\theta)-L(\theta)\right\}$
$\subseteq$ $\bigcup_{\theta,\theta^{\prime}\in\Theta}\left\{\hat{L}_{k_{2}}(\theta^{\prime})-L(\theta^{\prime})\leq-\max\left(\dfrac{\Delta_{o}}{2}-\dfrac{\delta}{4},\dfrac{\delta}{4}\right)\text{ or }\hat{L}_{k_{2}}(\theta)-L(\theta)\geq\max\left(\dfrac{\Delta_{o}}{2}-\dfrac{\delta}{4},\dfrac{\delta}{4}\right)\right\}$
$\subseteq$ $\bigcup_{\theta\in\Theta}\left\{\left|\hat{L}_{k_{2}}(\theta)-L(\theta)\right|\geq\max\left(\dfrac{\Delta_{o}}{2}-\dfrac{\delta}{4},\dfrac{\delta}{4}\right)\right\}$
where the last line holds because $\max\left(\Delta_{o}/2-\delta/4,\delta/4\right)\geq\Delta_{o}/4$. This gives
$T_{k_{2}}\left(\dfrac{\Delta_{o}}{4}\right)\leq\frac{1}{2\left|\Theta\right|}\text{\ \ for all $k_{2}\geq\underline{k}$}$
Due to the uniform convergence (49), we have
in probability, and hence
Finally, we combine all the pieces to get
$\left\{\hat{\theta}_{n}\not\in\Theta^{2\delta}\right\}$
@@
$\subseteq$ $\left\{\mathcal{S}\cap\Theta^{\delta}=\emptyset\right\}\cup\left\{\hat{\theta}_{n}\not\in\mathcal{S}^{\delta}\right\}$
By Lemma 6 we have $\mathbb{P}\left(\mathcal{S}\cap\Theta^{\delta}=\emptyset\right)\to 0$ under the conditions that $\limsup_{k\to\infty}\mathcal{E}_{k,\delta}<1$ and $k_{1},n/k_{1},B_{1}\to\infty$. Together with the condition $\mathbb{P}\left(\epsilon\geq\delta/2\right)\to 0$ and (50), we conclude $\mathbb{P}\left(\hat{\theta}_{n}\not\in\Theta^{2\delta}\right)\to 0$ by the union bound. $\Box$
Appendix D Supplementary material for numerical experiments

### v4

@@ Appendix D Proofs for Technical Lemmas
$\mathbb{P}\left(\mathbb{P}_{*}\left(\mathcal{A}(z_{1}^{*},\ldots,z_{k}^{*})\in\Theta^{\delta}\right)<\frac{1-\mathcal{E}_{k,\delta}}{e}\right)$$\leq$ $\min\left\{\exp\left(-\frac{n}{2k}\cdot\left(1-\frac{2}{e}\right)(1-\mathcal{E}_{k,\delta})\right),\left(\frac{e^{2}\mathcal{E}_{k,\delta}}{e-1}\right)^{(1-1/e)\frac{n}{2k}}\right\}.$
@@
$\mathbb{P}\left(\mathcal{S}\cap\Theta^{\delta}=\emptyset\right)\leq$ $\min\left\{\exp\left(-\frac{n}{2k}\cdot\left(1-\frac{2}{e}\right)(1-\mathcal{E}_{k,\delta})\right),\left(\frac{e^{2}\mathcal{E}_{k,\delta}}{e-1}\right)^{(1-1/e)\frac{n}{2k}}\right\}+\exp\left(-\frac{B_{1}(1-\mathcal{E}_{k,\delta})}{e}\right)$
$\leq$ $\min\left\{\exp\left(-\dfrac{1-2/e}{1-1/e}\cdot(1-\mathcal{E}_{k,\delta})\right),\frac{e^{2}\mathcal{E}_{k,\delta}}{e-1}\right\}^{(1-1/e)\frac{n}{2k}}+\exp\left(-\frac{B_{1}(1-\mathcal{E}_{k,\delta})}{e}\right).$
The conclusion then follows by setting $C_{4},C_{5},C_{6},C_{7}$ to be the appropriate constants. ∎
D.7 Proof of Lemma C.13
Let $\hat{L}_{k}(\theta):=\frac{1}{k}\sum_{i=1}^{k}l(\theta,z_{i})$. Let $\theta^{*}$ be an optimal solution of (1). We have
@@
This immediately gives (37). (38) is obvious given (36) and (37). ∎
Appendix E Linear Regression Example
In this section, we present another example demonstrating that our algorithm is capable of turning a polynomial-tailed base learner into exponential. We consider the linear regression problem
$\min_{\theta\in[-1,1]}\mathbb{E}\bigl{[}(x\theta-y)^{2}\bigr{]},$ (56)
where the data $\{(x_{i},y_{i})\}_{i=1}^{n}$ are i.i.d. samples such that $x_{i}\in\{-1,1\}$, and $y_{i}=x_{i}\theta^{*}+\epsilon_{i}$.
Assumption E.1.
The unknown true coefficient of problem (56) is $\theta^{*}=0$. The random variables $\{\epsilon_{i}\}_{i=1}^{n}$ are i.i.d. distributed with zero mean and symmetric with respect to 00. The second and forth moments of $\epsilon_{i}$ are finite, denoted as $\sigma^{2}=\mathbb{E}\left[\epsilon_{i}^{2}\right]$ and $\mu_{4}=\mathbb{E}\left[\epsilon_{i}^{4}\right]$. Moreover, there exist constants $C>0$ and $\alpha>0$ such that $P(\epsilon_{i}>t)>C(t+1)^{-\alpha}$ for all $t>0$, i.e., $\epsilon_{i}$ has a polynomial tail.
Under the setting described by (56) and Assumption E.1, the least-squares estimator of $\theta$ is given by
$\theta^{LS}_{n}=\mathcal{P}_{[-1,1]}\left(\frac{\sum_{i=1}^{n}x_{i}y_{i}}{\sum_{i=1}^{n}x_{i}^{2}}\right)=\mathcal{P}_{[-1,1]}\left(\frac{\sum_{i=1}^{n}x_{i}\epsilon_{i}}{\sum_{i=1}^{n}x_{i}^{2}}+\theta^{*}\right)=\mathcal{P}_{[-1,1]}\left(\frac{\sum_{i=1}^{n}x_{i}\epsilon_{i}}{n}\right),$
where $\mathcal{P}_{[-1,1]}(\cdot)$ denotes the projection operator onto the interval $[-1,1]$. Since the true coefficient $\theta^{*}=0$, for any estimator $\hat{\theta}$ that takes values between $[-1,1]$, its excess risk is equal to $(\hat{\theta})^{2}$. For instance, the excess risk of $\theta^{LS}_{n}$ can be expressed as
$\left(\theta^{LS}_{n}\right)^{2}=\min\left\{\left(\frac{\sum_{i=1}^{n}x_{i}\epsilon_{i}}{n}\right)^{2},1\right\}.$
Theorem E.2.
Under Assumption E.1, the followings hold true.
  • The excess risk of the least-squares estimator $\theta^{LS}_{n}$ exhibits a polynomial tail in $n$. Specifically, for every $\delta\in(0,1)$ and $n>1$, it holds that
    $\mathbb{P}\left((\theta^{LS}_{n})^{2}>\delta\right)>C(n\sqrt{\delta}+1)^{-\alpha}.$ (57)
  • Under our ensemble method, the excess risk of the output estimator $\hat{\theta}_{n}$ has an exponentially decreasing tail.
E.1 Proof of Theorem E.2
We first show the polynomial tail of excess risk for the least-squares estimator $\theta_{n}^{LS}$. For $k\leq n$, let $\bar{\epsilon}_{k}:=1/k\cdot{\sum_{i=1}^{k}\epsilon_{i}}$ be the sample average of the first $k$ noise terms. Then, it holds that $\bar{\epsilon}_{k+1}=\left(k\bar{\epsilon}_{k}+\epsilon_{k+1}\right)/(k+1)$.
For every $\delta\in(0,1)$ and $n>1$, we have that
@@ Appendix E Linear Regression Example
$\mathbb{P}\left((\theta^{LS}_{n})^{2}>\delta\right)=\mathbb{P}\left(\left(\frac{\sum_{i=1}^{n}x_{i}\epsilon_{i}}{n}\right)^{2}>\delta\right)=\mathbb{P}\left(\left(\frac{\sum_{i=1}^{n}\epsilon_{i}}{n}\right)^{2}>\delta\right)=2\mathbb{P}\big{(}\bar{\epsilon}_{n}>\sqrt{\delta}\big{)},$ (58)
where we used the symmetry of $\epsilon_{i}$ and that $x_{i}\in\{-1,1\}$. Then, using the recursive relation between $\bar{\epsilon}_{n}$ and $\bar{\epsilon}_{n-1}$, we can further show that
@@
$\mathbb{P}\big{(}\bar{\epsilon}_{n}>\sqrt{\delta}\big{)}$ $\geq\mathbb{P}\left(\frac{(n-1)\bar{\epsilon}_{n-1}}{n}\geq 0\text{ and }\frac{\epsilon_{n}}{n}>\sqrt{\delta}\right)$ (59)
@@
$=\mathbb{P}\left(\frac{(n-1)\bar{\epsilon}_{n-1}}{n}\geq 0\right)\cdot\mathbb{P}\left(\frac{\epsilon_{n}}{n}>\sqrt{\delta}\right)$
$>\mathbb{P}\left(\bar{\epsilon}_{n-1}\geq 0\right)\cdot C(n\sqrt{\delta}+1)^{-\alpha},$
where the second line is due to the independence between $\bar{\epsilon}_{n-1}$ and $\epsilon_{n}$ in the second line, and the last line uses Assumption E.1. Since each $\epsilon_{i}$ is symmetric, we have $\mathbb{P}\left(\bar{\epsilon}_{n-1}\geq 0\right)=1/2$. Hence, the proof of (57) is completed by combining (58) and (59).
Now, we proceed to show the exponential tail of excess risk for our ensemble method (Algorithm 2), where the proof is based on Theorem C.11, i.e., the formal version of Theorem 2.3. To apply the bound (34) from Theorem C.11, we need to derive upper bounds for the following two quantities: the empirical process tail $T_{k}(\cdot)$, and the excess risk tail of the base learner, i.e., the least-squares estimator $\theta^{LS}_{k}$ with $k$ samples.
For any $t>0$, we can show that the empirical process tail satisfies that
$T_{k}(t)$ $=\mathbb{P}\left(\sup_{\theta\in[-1,1]}\left|\frac{1}{k}\sum_{i=1}^{k}(x_{i}\theta-y_{i})^{2}-\mathbb{E}[(x\theta-y)^{2}]\right|>t\right)$ (60)
$=\mathbb{P}\left(\sup_{\theta\in[-1,1]}\left|\frac{1}{k}\sum_{i=1}^{k}(x_{i}\theta-\epsilon_{i})^{2}-(\theta^{2}+\sigma^{2})\right|>t\right)$
@@
$=\mathbb{P}\left(\sup_{\theta\in[-1,1]}\left|\frac{1}{k}\sum_{i=1}^{k}\epsilon_{i}^{2}-\sigma^{2}-2\theta\cdot\frac{1}{k}\sum_{i=1}^{k}x_{i}\epsilon_{i}\right|>t\right)$
$\leq\mathbb{P}\left(\left|-\sigma^{2}+\frac{1}{k}\sum_{i=1}^{k}\epsilon_{i}^{2}-\frac{2}{k}\sum_{i=1}^{k}x_{i}\epsilon_{i}\right|>t\right)+\mathbb{P}\left(\left|-\sigma^{2}+\frac{1}{k}\sum_{i=1}^{k}\epsilon_{i}^{2}+\frac{2}{k}\sum_{i=1}^{k}x_{i}\epsilon_{i}\right|>t\right),$
where the last inequality uses the union bound and the observation that the maximum of the absolute value term is achieved either at $\theta=1$ or $\theta=-1$. Using the symmetry of $\epsilon_{i}$ and the fact that $x_{i}\in\left\{-1,1\right\}$, we further derive from (60) that
$T_{k}(t)$ $\leq 2\mathbb{P}\left(\left|\frac{1}{k}\sum_{i=1}^{k}\epsilon_{i}^{2}-\sigma^{2}-\frac{2}{k}\sum_{i=1}^{k}\epsilon_{i}\right|>t\right)$ (61)
$\leq 2\mathbb{P}\left(\left|\frac{1}{k}\sum_{i=1}^{k}\epsilon_{i}^{2}-\sigma^{2}\right|>\frac{t}{2}\right)+2\mathbb{P}\left(\left|\frac{2}{k}\sum_{i=1}^{k}\epsilon_{i}\right|>\frac{t}{2}\right)$
$\leq\frac{8\mu_{4}}{kt^{2}}+\frac{32\sigma^{2}}{kt^{2}},$
where we apply the union bound again in the second line and use the Markov’s inequality in the last line.
Now, we assess the excess risk tail of the least-squares estimator $\theta^{LS}_{k}$. For $\delta<1$, similar as (58), we can apply the Markov’s inequality to show that
@@
$\mathcal{E}_{k,\delta}=\mathbb{P}\left((\theta^{LS}_{k})^{2}>\delta\right)=\mathbb{P}\left(\left(\frac{\sum_{i=1}^{k}\epsilon_{i}}{k}\right)^{2}>\delta\right)\leq\frac{\sigma^{2}}{k\delta}.$ (62)
By instantiating (34) in Theorem C.11 with the tail bounds on $T_{k}(t)$ and $\mathcal{E}_{k,\delta}$ given by (61) and (62), we can finally obtain the following tail bound on the excess risk of our estimator $\hat{\theta}_{n}$:
$\mathbb{P}\left((\hat{\theta}_{n})^{2}>\delta\right)$ $\leq B_{1}\left(3\min\left\{e^{-2/5},C_{1}\frac{32(\mu_{4}+4\sigma^{2})}{k_{2}\min\left\{\underline{\epsilon},\delta-\overline{\epsilon}\right\}^{2}}\right\}^{\frac{n}{2C_{2}k_{2}}}+e^{-B_{2}/C_{3}}\right)$ (63)
$\quad\ +\min\left\{e^{-\big{(}1-\frac{\sigma^{2}}{k_{1}\delta}\big{)}\big{/}C_{4}},C_{5}\frac{\sigma^{2}}{k_{1}\delta}\right\}^{\frac{n}{2C_{6}k_{1}}}+e^{-B_{1}\big{(}1-\frac{\sigma^{2}}{k_{1}\delta}\big{)}\big{/}C_{7}},$
for every $k_{1},k_{2}\leq n$ and $\delta\in(0,1)$ such that $\delta>\overline{\epsilon}$, $\frac{32(\mu_{4}+4\sigma^{2})}{k_{2}(\delta-\overline{\epsilon})^{2}}+\frac{32(\mu_{4}+4\sigma^{2})}{k_{2}\underline{\epsilon}^{2}}<\frac{1}{5}$, and $\frac{\sigma^{2}}{k_{1}\delta}<1$. Note that these conditions guarantee that the upper bound in (63) is meaningful and that $T_{k_{2}}((\delta-\overline{\epsilon})/2)+T_{k_{2}}(\underline{\epsilon}/2)<1/5$, as required by Theorem 2.3. Therefore, we conclude that the excess risk for the output solution $\hat{\theta}_{n}$ of my ensemble method has an exponential tail, which completes the proof of Theorem E.2. ∎
Appendix F Supplementary Material for Numerical Experiments